factorize
2x^2+2(squareroot)6xy + 3y^2
1
Expert's answer
2012-06-19T10:39:38-0400
we use (a+b)^2=a^2+2ab+b^2
2x^2+2(squareroot)6xy +
3y^2=(sqrt(2)*x)^2+2*sqrt(2)*x*sqrt(3)*y+(sqrt(3)*y)^2=use formula=(
sqrt(2)*x+sqrt(3)*y )^2